Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Coin Flip Strategy

Expert replies
by priyankamishra11 » Sat Sep 27, 2008 4:08 pm
Hi ,

I wanted to apply coin flip strategy given by Stuart " www.beatthegmat.com/coin-flip-questions ... 17911.html " in the following question, but I am not getting the answer can someone please explain what is wrong?

If a certain coin is flipped, the probability that the coin will land heads is 1/2. If the coin is flipped 5 times, what is the probability that it will land heads up on the first 3 flips and not on the last 2 flips?

A. 3/5
B. 1/2
C. 1/5
D. 1/8
E. 1/32
Regards,
Priyanka
Join the discussion
Source: — Problem Solving |

by cramya » Sat Sep 27, 2008 4:19 pm
I am also eager to know how to apply the pascal triangle coin flip strategy here.



I just did

p(hhhtt) = 1/2*1/2*1/2*1/2*1/2 = 1/32
Join the discussion

by priyankamishra11 » Sat Sep 27, 2008 4:33 pm
yes thats how I did it, but I am not getting this same answer by applying either coin flip formula or pascal triangle.

Can someone explain it?
Regards,
Priyanka
Join the discussion

by stop@800 » Sun Sep 28, 2008 6:30 am
Total possible outcomes = 2^5 = 32
Favorable is just one ie HHHTT
so
Ans IMO is 1/32
Join the discussion

by stop@800 » Sun Sep 28, 2008 6:37 am
You need 3 heads out of 3 and
2 tails out of 2

We need to carefully apply pascal triangle.

We need 3 heads out of first 3 toss so apply pascal of 3 coins
and 2 tails from last 2 tails so apply pascal of 2 coins
and multiple those

from 1331
we will get 1 favorable and total 8

for two coins it is
121
hence 1 / 4

so 1/4 * 1/8

= 1 / 32

Hope this helps!!
Join the discussion

Re: Coin Flip Strategy

by Stuart@KaplanGMAT » Sun Sep 28, 2008 7:59 pm
priyankamishra11 wrote:Hi ,

I wanted to apply coin flip strategy given by Stuart " www.beatthegmat.com/coin-flip-questions ... 17911.html " in the following question, but I am not getting the answer can someone please explain what is wrong?

If a certain coin is flipped, the probability that the coin will land heads is 1/2. If the coin is flipped 5 times, what is the probability that it will land heads up on the first 3 flips and not on the last 2 flips?

A. 3/5
B. 1/2
C. 1/5
D. 1/8
E. 1/32
This isn't a great question for the Triangle - there's a much easier way to solve.

We know that probability = #desired/total#

In this case, we have only 1 desired outcome: HHHTT

We know that for 5 flips, there are 2^5 = 32 total possible outcomes.

So, the answer is simply 1/32.

I suppsed you could make things far more complicated as stop@800 did and apply the Triangle that way (his solution is perfect), but that seems like a lot of extra effort!
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by cramya » Mon Sep 29, 2008 7:05 pm
Thanks Stuart!
Join the discussion

by walker » Thu Apr 09, 2009 9:37 am
Agreed with stuart

probability = #desired/total#
Join the discussion